OpenStudy (anonymous):

Raina runs 7 miles in 60 minutes. At the same rate, how many miles would she run in 24 minutes?

OpenStudy (anonymous):

Raina runs 7 miles in 60 minutes. So we have the proportion of \[\frac{7 \text{ miles }}{60 \text{ minutes }}\]

OpenStudy (anonymous):

Now using proprtions we determine how much Raina will run in 24 minutes \[\frac{7 \text{ miles }}{60 \text{ minutes }}=\frac{x\text{ miles }}{24 \text{ minutes }}\]

OpenStudy (anonymous):

Now we cross multiply and get \[7*24=60x\] \[168=60x\] \[\frac{168}{60}=x\] \[x=2.8\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Do you follow?? Any questions?